A parallel plate air capacitor has a capacitance of 5 F. What is the capacitance if the separation between the plates is decreased to one-half its original value while keeping all other factors constant?

1
Expert's answer
2022-03-21T14:06:53-0400

The capacitance of parallel plate capacitor

"C=\frac{\epsilon_0 A}{d}"

"C'=\frac{\epsilon_0 A}{d'}=\frac{\epsilon_0 A}{d/2}=2\times\frac{\epsilon_0 A}{d}=2C\\
=2\times \rm 5\: F=10\: F"